OpenStudy (anonymous):

What's the derivative of ln(e^(dx)+k), assuming that d and k are constants. Thanks!

OpenStudy (anonymous):

Given something of the form \(\ln(\text{inside})\), its derivative is \(\dfrac{(\text{inside})'}{(\text{inside})}\). Here, \(\text{inside}=e^{dx}+k\).

OpenStudy (anonymous):

So would it be ((e^(dx)+1)/(e^(dx)+k)?

hartnn (hartnn):

your denominator is correct, but what about the derivative of inside = e^dx+k ??

OpenStudy (anonymous):

Don't I differentiate the derivative of the inside, so derivative of e^dx and the derivative of k. Derivatives of e is itself isn't it? And since k is a constant, it'd be 0?

hartnn (hartnn):

yes, derivative of k is 0, (you have written 1 there instead) and derivative of e^x is e^x, yes. but derivative of e^dx will be d*e^dx because we multiply by the co-efficient of x, because of chain rule. d/dx (e^dx ) = e^dx * d/dx (dx) = e^dx *d understand ?

OpenStudy (anonymous):

Yeah I just realized that as I was replying, sorry. But thanks a lot! I get it now :)

hartnn (hartnn):

welcome ^_^ and hey, since you are new here. \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{ }}\color{#e6e600}{\text{S}}\color{green}{\text{t}}\color{blue}{\text{u}}\color{purple}{\text{d}}\color{purple}{\text{y}}\color{red}{\text{!}}\color{orange}{\text{!}}\color{#e6e600}{\text{ }}\color{green}{\text{:}}\color{blue}{\text{)}}\color{purple}{\text{}}\end{array} \]

OpenStudy (anonymous):

Thanks! This site is very helpful, and I'm liking the Goku picture!

hartnn (hartnn):

hehe, he is my favorite superhero :D